OpenStudy (anonymous):

pls. . . help me. . The sides of a parallelogram are 55 cm and 71 cm. Find the length of each diagonal if the longer angle measures 106 degrees. 2.a sample of 3000 observations has a mean of 82 and a standard deviation of 16. Using the empirical rule, find the percentage of the observations fall in the intervals x=+-25. 3.using the above data, how many samples do not belong to +-25?

OpenStudy (ash2326):

@kris10 Do you know sin formula for triangles?

OpenStudy (ash2326):

I mean sine oops

OpenStudy (anonymous):

opposite/hypotenuse

OpenStudy (campbell_st):

so you can use the cosine rule with a = 55, b = 71 and C = 106 so let c be the longer diagonal \[c^2 = 55^2 + 71^2 - 2 \times 55 \times 71 \times \cos(106)\] solve for c let the shorter diagonal be d the the angle will be 74 so the method is identical \[d^2 = a^2 + b^2 - 2ab \cos(D)\]

OpenStudy (campbell_st):

oops I forgot the diagram  wich gives 2 triangles 

OpenStudy (anonymous):

im confused